OpenStudy (anonymous):

solve p=2(l+w) for l

OpenStudy (anonymous):

So this is a perimeter problem?

OpenStudy (anonymous):

i know the answer is l=p/2 - w i just want to know how to do it

OpenStudy (anonymous):

first distribute. p=2l+2w then subtract 2w Now your left with p-2w=2l divide by 2 and your final answer is l=p/2-w

OpenStudy (anonymous):

hope this helps =]